# Group-aware Parameter-efficient Updating for Content-Adaptive Neural Video Compression

### v1

@@ 2.3. Parameter-Efficient Transfer Learning
Our PETL strategy is closely aligned with the concept of Low-Rank Adaptation (He et al., 2021; Zhu et al., 2021; Hu et al., 2021b; He et al., 2022; Pfeiffer et al., 2020; Karimi Mahabadi et al., 2021; Edalati et al., 2022), which are recognized for their practicality and efficiency in delta-tuning methods by using adaptors. Various modifications to adaptors have been explored. These include adjustments in the placement of adaptors within the model (He et al., 2021; Zhu et al., 2021), the application of targeted pruning techniques to enhance model efficiencys (He et al., 2022), and the adoption of reparametrization strategies to optimize performance(Karimi Mahabadi et al., 2021; Edalati et al., 2022). Such innovations demonstrate the adaptability and versatility of adaptors in enhancing task-specific model performance.
To the best of our knowledge, our work represents the pioneering effort to integrate the PETL strategy within the NVC framework. Specifically, we introduce two distinct adaptor configurations: parallel and series adaptors. The parallel adaptor is designed for use with large-parameter modules, such as the motion and contexture feature encoding network. Conversely, the series adaptor is tailored for light-weight modules, including motion estimation, motion encoding, motion hyper-prior encoding and contexture hyper-prior encoding modules. These dual-adaptor styles significantly enhance our ability to optimize the NVC process during encoding.
Figure 1. The overview of our Patch-based Groups of Pictures (GoP) updating strategy. It begins by segmenting each video into multiple GoPs as in conventional video codecs. We then spatially divide each GoP into several patch-based GoPs. Following this, we apply our Group-aware Parameter-efficient Updating (GPU) technique to update these patch-based GoPs sequentially. Once all the patch-based GoPs have been optimized, we feed the integrated GoP into the Neural Video Compression (NVC) framework to generate the compressed bit-stream. This process is repeated iteratively until the entire video has been compressed.
@@ 3.1.1. Patch-based GoP Updating
To compress a specified video sequence $\mathcal{X}=\{X_{1},X_{2},\dots,X_{t-1},X_{t},\dots\}$, we initially divide it into $K$ non-overlap groups of pictures (GoP), each consisting of $T$ frames, denoted as $\mathcal{G}_{k}=\{X_{t},X_{t+1},\dots,X_{t+T-1}\}$. This process results in $\mathcal{X}$ being represented as $\mathcal{X}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{K}\}$. In this work, we adopt the widely used predictive-frame (P-Frame) coding approach in NVC, as referenced in (Hu, 2020; Hu et al., 2021a; Li et al., 2021, 2022, 2023; Chen et al., 2023; Hu et al., 2022; Zhu et al., 2022). We designate the first frame in each GoP as an intra-coding frame (I-Frame) and the subsequent frames as P-Frames, following a pattern of $\{IPPPP\cdots\}$ and only apply our updating mechanism upon P-frames.
@@
Our goal is to perform our online updating mechanism, GPU, for each GoP $\mathcal{G}_{k}$ to minimize error propagation. However, directly updating $\mathcal{G}_{k}$ poses significant challenges due to the potentially high resolution and large number of frames involved, leading to substantial memory and computational demands. To address this, we spatially partition $\mathcal{G}_{k}$ into $N$ patch-based GoPs $\mathcal{G}_{k}^{i}$, each located at a specific spatial position $i$. This allows us to apply the updating mechanism on each $\mathcal{G}_{k}^{i}$ individually on the GPU, using the objective function $\mathcal{L}$, which will be elaborated in Section 3.2.2. After updating all patch-based GoPs $\mathcal{G}_{k}^{i}$, we proceed to compress the entire GoP $\mathcal{G}_{k}$ into a bit-stream utilizing our NVC framework with the newly updated parameters on the encoder side. This process is iteratively carried out for all $\mathcal{G}_{k}$s. We demonstrate this algorithm 1 and its implementation details in Fig. 1.
@@ 3.1.2. Encoder-side Adaptor
Existing content-adaptive NVC (Lu et al., 2019; Tang et al., 2023) methods often update all parameters in the encoding modules directly. While this full-updating strategy may naturally improve performance, it also significantly increases computational costs. Such a strategy becomes impractical for modern NVC systems (Li et al., 2023), which typically involve a vast number of parameters.
To enhance our updating process’s efficiency, we have implemented the adaptor mechanism, a technique widely used in PETL methods (Hu et al., 2021b). As shown in Fig. 2 (c), our approach $\delta(\cdot)$ utilizes a streamlined architecture comprising merely three convolutional layers $\mathbf{W}$ and two LeakyReLU activation layers $\sigma(\cdot)$. This structure is designed to process input features, denoted as $\mathbf{F}_{in}\in\mathcal{R}^{C_{in}\times H_{in}\times W_{in}}$ and generate updated features $\mathbf{F}\in\mathcal{R}^{C\times H\times W}$ using adaptors. Initially, we apply a $1\times 1$ convolutional layer for pre-processing, characterized by weights $\mathbf{W}_{pre}\in\mathcal{R}^{C_{in}\times C_{dw}\times 1\times 1}$ to reshape the feature into a channel-squeezed form $\mathbf{F}\in\mathcal{R}^{C_{dw}\times H\times W}$ s.t., $C_{dw}<C_{in}$ $C_{dw}<C$. Following the strategy commonly employed in Low-Rank Adaptation methods (Hu et al., 2021b; Shen et al., 2023), we then apply a depth-wise convolution, with weights $\mathbf{W}_{dw}\in\mathcal{R}^{C_{dw}\times M\times M}$ and a kernel size of $M$ to produce the feature $\mathbf{F}\in\mathcal{R}^{C_{dw}\times H\times W}$. Finally, we employ a $1\times 1$ (i.e., point-wise) convolutional layer with zero-initialized weights, $\mathbf{W}_{zero}\in\mathcal{R}^{C_{dw}\times C\times 1\times 1}$ inspired by (Zhang et al., 2023). Initially, this results in the delta feature $\delta(\mathbf{F}_{in})$ having no impact on the feature $\mathbf{F}_{in}$. Over the updating process, the impact of $\delta(\mathbf{F}_{in})$ will incrementally increase from zero, optimizing the parameters to enhance the overall model performance. The entire updating can be concisely represented by the following equation, where $\times$ and $\otimes$ denote the convolutional and depth-wise convolutional operations:
@@
3.2. Content Adaptive NVC with Adaptors
@@ 3.2.1. Overview
We deploy our proposed adaptors in an NVC framework as shown in Fig. 3. Our framework compresses the current frame $X_{t}$ to obtain the reconstructed frame $\hat{X}_{t}$ and associated quantized features $\hat{\mathbf{M}}_{t}$ and $\hat{\mathbf{Y}}_{t}$, where the subscript $t$ represents the current time-step $t$. The process involves four main steps:  1. Motion Estimation and Compression;  2. Motion Reconstruction and Compensation;  3. Contexture Feature Compression; and  4. Frame Reconstruction. In general, we follow the previous content-adaptive NVC methods (Lu et al., 2020b; Tang et al., 2023) and only update the modules, that can only be used during the encoding procedure (i.e., in Steps 1 & 3) and leave those modules, which are used in decoding stages (i.e., in Step 2 & Step 4) unaltered. The details are given as follows:
@@
Step 1. Motion Estimation and Compression. We first estimate the raw motion $V_{t}$ between reference frame (i.e., previously reconstructed frame) $\hat{X}_{t-1}$ and current-coding frame $X_{t}$ through the Motion Estimation operation, where we directly adopt a compact optical flow network SpyNet (Ranjan and Black, 2017) as in (Lu et al., 2019; Li et al., 2021, 2022, 2023; Chen et al., 2023). Following this, we employ an auto-encoder-style module, Motion Encoder, to compress $V_{t}$ into the motion feature $\mathbf{M}_{t}$, followed by quantizing it to $\hat{\mathbf{M}}_{t}$. To losslessly transmit $\hat{\mathbf{M}}_{t}$, we need to adopt arithmetic encoder (resp., decoder) to losslessly encode to (resp., decode from) the bit-stream. To minimize the bit-rates, we employ a hyper-prior-based entropy model as in (Li et al., 2023), which enhances the estimation of $\hat{\mathbf{M}}_{t}$’s distribution and improves the efficiency of cross-entropy coding. Consequently, this phase involves generating a quantized motion hyper-prior feature by using a Motion Hyper-Prior Encoder. In this step, we can update the Motion Estimation, Motion Encoder and Motion Hyper-Prior Encoder online during the encoding procedure, aiming to compress the quantized motion feature $\hat{\mathbf{M}}_{t}$ in a better way. Specifically, we incorporate serial adaptors for those light-weight coding modules, Motion Estimation and Motion Hyper-Prior Encoder, while for more complex coding modules, Motion Encoder, parallel adaptors are employed to facilitate improvements.
@@
Step 2. Motion Reconstruction and Compensation. Using the transmitted quantized motion feature $\hat{\mathbf{M}}_{t}$, we can proceed with motion compensation operation. This process begins with the reconstruction of $\hat{\mathbf{M}}_{t}$ into the reconstructed motion $\hat{V}_{t}$ by employing Motion Reconstruction modules that consist of several standard convolutional blocks. Once we have $\hat{V}_{t}$, the motion compensation procedure is carried out to generate warped features $\bar{\mathbf{F}}_{t}$, effectively reducing temporal redundancy. In this work, we apply a multi-scale temporal context extraction strategy, directly following (Sheng et al., 2022) to obtain multiple scales of temporal context information $\bar{\mathbf{F}}_{t}$. Since the modules in this phase are integral to both the encoding and decoding stages, we do not update them during the encoding process.
@@
Step 3. Contexture Feature Compression. Adhering to the methodologies established by current deep contextual coding frameworks (Li et al., 2022; Sheng et al., 2022; Li et al., 2023; Chen et al., 2023), we employ a multi-scale spatial-temporal Contexture Feature Encoder. This encoder generates the latent feature $\mathbf{Y}_{t}$ using the current frame $X_{t}$ and previously generated warped features $\bar{\mathbf{F}}_{t}$. Similar to Step 1, we again quantize $\mathbf{Y}_{t}$ into $\hat{\mathbf{Y}}_{t}$ and subsequently transmit it losslessly using arithmetic coding alongside a hyper-prior-based entropy model. Consequently, a Contexture Hyper-Prior Feature Encoder is deployed to produce a hyper-prior feature. In this step, we enable the online updating of the Contexture Feature Encoder and the Contexture Hyper-Prior Feature Encoder. For the more complex Contexture Feature Encoder, parallel adaptors are integrated, whereas for the more compact Contexture Hyper-Prior Feature Encoder, serial adaptors are utilized.
@@
Step 4. Frame Reconstruction. Finally, we revert the quantized latent feature $\hat{\mathbf{Y}}_{t}$ back into the reconstructed frame $\hat{X}_{t}$ by utilizing a multi-scale spatial-temporal Contexture Feature Decoder, with assistance from the warped features $\bar{\mathbf{F}}_{t}$. This step is only used during the decoding process, hence there is no updating of any modules.
@@ 4. Experiments
Figure 4. Rate-distortion (i.e., bit-rates vs PSNR) performance comparison between our and other state-of-the-art video compression methods on the standard video compression benchmarks, HEVC Class B, C, D, and E.
@@ 4.1.1. Datasets
To evaluate the effectiveness of our approach, we implemented it on video sequences from four of the most representative video compression datasets: HEVC (Sullivan et al., 2012) Class B, C, D, and E. These datasets are selected by the standardization committees ISO/IEC and ITU-T and are extensively utilized in the assessment of new video coding technologies. They encompass videos with resolutions ranging from 416$\times$240 to 1920$\times$1080 featuring a diverse array of video scenes. Furthermore, we assessed the adaptability of our method using the most widely used cardiac medical MRI sequence dataset, ACDC (Bernard et al., 2018), where we utilized 10 scans from Patient 101, each with a resolution of $256\times 232$. We measured the bit-rates by bits-per-pixel (bpp) and assessed the quality of compression by using the peak signal-to-noise ratio (PSNR).
@@ 4.1.2. Baseline Codecs
Our evaluation aimed to assess the efficacy of our method relative to both traditional video codecs and a range of NVC methods, including standard and content-adaptive variants. For traditional codecs, we utilized the H.265/HEVC (Sullivan et al., 2012) and H.266/VVC (Bross et al., 2021) standards, employing their reference software versions HM-16.20 (HM, [n. d.]) and VTM-11.2 (vtm, [n. d.]), respectively. Our analysis also included two leading NVCs as baselines: DCVC (Li et al., 2021) and its enhanced version, DCVC_DC (Li et al., 2023), with the latter representing the latest advancement in NVC study to the best of our knowledge. It surpasses traditional video codecs in performance, sharing the high-level syntax with DCVC but featuring a more intricate architecture. Additionally, we examined two content-adaptive NVC methods: one by Lu et al. (Lu et al., 2020b) and another by Tang et al. (Tang et al., 2023) with Tang’s method being the most recent development in the field of content-adaptive NVC to the best of our knowledge. To ensure a fair comparison, we recalibrated the performance metrics for DCVC, DCVC_DC, H.266, and H.265 using publicly available codebases and pre-trained models. Our setup involved an I-Frame interval of 32 and the encoding of 96 frames from each video, consistent with (Li et al., 2023). For each MRI scan from the ACDC dataset, we adopted an I-Frame interval of 30 and encoded 30 slices. For the other codecs, due to the unavailability of public implementations, we relied on performance data as reported in their original studies.

### v2

@@ 2.3. Parameter-Efficient Transfer Learning
Our PETL strategy is closely aligned with the concept of Low-Rank Adaptation (He et al., 2021; Zhu et al., 2021; Hu et al., 2021b; He et al., 2022; Pfeiffer et al., 2020; Karimi Mahabadi et al., 2021; Edalati et al., 2022), which are recognized for their practicality and efficiency in delta-tuning methods by using adaptors. Various modifications to adaptors have been explored. These include adjustments in the placement of adaptors within the model (He et al., 2021; Zhu et al., 2021), the application of targeted pruning techniques to further enhance model efficiencys (He et al., 2022), and the adoption of reparametrization strategies to optimize performance(Karimi Mahabadi et al., 2021; Edalati et al., 2022). Such innovations demonstrate the adaptability and versatility of adaptors in enhancing task-specific model performance.
To the best of our knowledge, our work represents the pioneering effort to integrate the PETL strategy within the NVC framework. Specifically, we introduce two distinct adaptor configurations: parallel and series adaptors. The parallel adaptor is designed for use with large-parameter modules, such as the motion and contexture feature encoding networks. Conversely, the series adaptor is tailored for light-weight modules, including motion estimation, motion hyper-prior encoding and contexture hyper-prior encoding modules. Such a dual-configuration style significantly enhances our ability to optimize the NVC process during encoding.
Figure 1. The overview of our Patch-based Group of Pictures (GoP) updating strategy. It begins by segmenting each video into multiple GoPs as in conventional video codecs. We then spatially divide each GoP into several patch-based GoPs. Following this, we apply our Group-aware Parameter-efficient Updating (GPU) technique to update these patch-based GoPs sequentially. Once all the patch-based GoPs have been optimized, we feed the integrated GoP into the Neural Video Compression (NVC) framework to generate the compressed bit-stream. This process is repeated iteratively until the entire video has been compressed.
@@ 3.1.1. Patch-based GoP Updating
To compress a specified video sequence $\mathcal{X}=\{X_{1},X_{2},\dots,X_{t-1},X_{t},\dots\}$, we initially divide it into $K$ non-overlap GoPs, each consisting of $T$ frames, denoted as $\mathcal{G}_{k}=\{X_{t},X_{t+1},\dots,X_{t+T-1}\}$. This process results in $\mathcal{X}$ being represented as $\mathcal{X}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{K}\}$. In this work, we adopt the widely used predictive-frame (P-Frame) coding approach in NVC, as referenced in (Lu et al., 2019; Hu et al., 2021a; Li et al., 2021, 2022, 2023; Chen et al., 2023; Hu et al., 2022; Zhu et al., 2022). We designate the first frame in each GoP as an intra-coding frame (I-Frame) and the subsequent frames as P-Frames, following a pattern of $\{IPPPP\cdots\}$ and only apply our updating mechanism upon P-frames.
@@
Our goal is to perform our online updating mechanism, GPU, for each GoP $\mathcal{G}_{k}$ to minimize error propagation. However, directly updating $\mathcal{G}_{k}$ poses significant challenges due to the potentially high resolution and large number of frames involved, leading to substantial optimizing memory and computational demands. To address this, we spatially partition $\mathcal{G}_{k}$ into $N$ patch-based GoPs $\mathcal{G}_{k}^{i}$, each located at a specific spatial position $i$. This allows us to apply the updating mechanism on each $\mathcal{G}_{k}^{i}$ individually using the objective function $\mathcal{L}$, which will be elaborated in Section 3.2.2. After updating all patch-based GoPs $\mathcal{G}_{k}^{i}$, we proceed to compress the entire GoP $\mathcal{G}_{k}$ into a bit-stream utilizing our NVC framework with the newly updated parameters on the encoder side. This process is iteratively carried out for all GoPs $\mathcal{G}_{k}$. We demonstrate this algorithm 1 and its implementation details in Fig. 1.
@@ 3.1.2. Encoder-side Adaptor
Existing content-adaptive NVC (Lu et al., 2020; Tang et al., 2023) methods often update all parameters in the encoding modules directly. While this full-updating strategy may naturally improve performance, it also significantly increases computational costs. Such a strategy becomes impractical for modern NVC systems (Li et al., 2023), which typically involve a vast number of parameters.
To enhance our updating process’s efficiency, we have implemented the adaptor mechanism, a technique widely used in PETL methods (Hu et al., 2021b). As shown in Fig. 2 (c), our approach $\delta(\cdot)$ utilizes a streamlined architecture comprising merely three convolutional layers $\mathbf{W}$ and two LeakyReLU activation layers $\sigma(\cdot)$. This structure is designed to process input features, denoted as $\mathbf{F}_{in}\in\mathcal{R}^{C_{in}\times H_{in}\times W_{in}}$ and generate updated features $\mathbf{F}\in\mathcal{R}^{C\times H\times W}$ using adaptors. Initially, we apply a $1\times 1$ convolutional layer for pre-processing, characterized by weights $\mathbf{W}_{pre}\in\mathcal{R}^{C_{in}\times C_{dw}\times 1\times 1}$ to reshape the feature into a channel-squeezed form $\mathbf{F}\in\mathcal{R}^{C_{dw}\times H\times W}$. Following the strategy commonly employed in Low-Rank Adaptation methods (Hu et al., 2021b; Shen et al., 2023), we then apply a depth-wise convolution operation, with weights $\mathbf{W}_{dw}\in\mathcal{R}^{C_{dw}\times M\times M}$ and a kernel size of $M$ to produce the feature $\mathbf{F}\in\mathcal{R}^{C_{dw}\times H\times W}$. Finally, we employ a $1\times 1$ (i.e., point-wise) convolutional layer with zero-initialized weights, $\mathbf{W}_{zero}\in\mathcal{R}^{C_{dw}\times C\times 1\times 1}$ inspired by (Zhang et al., 2023). Initially, this results in the delta feature $\delta(\mathbf{F}_{in})$ having no impact on the feature $\mathbf{F}_{in}$. Over the updating process, the impact of $\delta(\mathbf{F}_{in})$ will incrementally increase from zero, optimizing the parameters to enhance the overall compression performance. The entire updating can be concisely represented by the following equation, where $\times$ and $\otimes$ denote the convolutional and depth-wise convolutional operations:
@@
3.2. Content-Adaptive NVC with Adaptors
@@ 3.2.1. Overview
We deploy our proposed adaptors in an NVC framework as shown in Fig. 3. Our framework compresses the current frame $X_{t}$ to obtain the reconstructed frame $\hat{X}_{t}$ and associated quantized features $\hat{\mathbf{M}}_{t}$ and $\hat{\mathbf{Y}}_{t}$, where the subscript $t$ represents the current time-step $t$. The process involves four main steps:  1. Motion Estimation and Compression;  2. Motion Reconstruction and Compensation;  3. Contexture Feature Compression; and  4. Frame Reconstruction. In general, we follow the previous content-adaptive NVC methods (Lu et al., 2019; Tang et al., 2023) and only update the modules, which can only be used during the encoding procedure (i.e., in Steps 1 & 3) and leave those modules, which are also used in decoding stages (i.e., in Step 2 & Step 4) unaltered. The details are as follows:
@@
Step 1. Motion Estimation and Compression. We first estimate the raw motion $V_{t}$ between reference frame (i.e., previously reconstructed frame) $\hat{X}_{t-1}$ and current-coding frame $X_{t}$ through the Motion Estimation operation, where we directly adopt a compact optical flow network SpyNet (Ranjan and Black, 2017) as in (Lu et al., 2019; Li et al., 2021, 2022, 2023; Chen et al., 2023). Following this, we employ an auto-encoder-style module, Motion Encoder, to compress $V_{t}$ into the motion feature $\mathbf{M}_{t}$, followed by quantizing it to $\hat{\mathbf{M}}_{t}$. To losslessly transmit $\hat{\mathbf{M}}_{t}$, we need to adopt arithmetic encoder (resp., decoder) to losslessly encode to (resp., decode from) the bit-stream. To minimize the bit-rates, we employ a hyper-prior-based entropy model as in (Li et al., 2023), which enhances the estimation of $\hat{\mathbf{M}}_{t}$’s distribution and improves the efficiency of cross-entropy coding. Consequently, this phase involves generating a quantized motion hyper-prior feature by using a Motion Hyper-prior Encoder. In this step, we can update the Motion Estimation, Motion Encoder and Motion Hyper-prior Encoder online during the encoding procedure, aiming to compress the quantized motion feature $\hat{\mathbf{M}}_{t}$ in a better way. Specifically, we incorporate serial adaptors for those light-weight coding modules, Motion Estimation and Motion Hyper-prior Encoder, while for more complex coding module, Motion Encoder, parallel adaptors are employed to facilitate improvements.
@@
Step 2. Motion Reconstruction and Compensation. Using the transmitted quantized motion feature $\hat{\mathbf{M}}_{t}$, we can proceed with motion compensation operation. This process begins with the decoding of $\hat{\mathbf{M}}_{t}$ into the reconstructed motion $\hat{V}_{t}$ by employing Motion Reconstruction modules that consist of several standard convolutional blocks. Once we have $\hat{V}_{t}$, the motion compensation procedure is carried out to generate warped features $\bar{\mathbf{F}}_{t}$, effectively reducing temporal redundancy. In this work, we apply a multi-scale temporal context extraction strategy, directly following (Sheng et al., 2022) to obtain multiple scales of temporal context information $\bar{\mathbf{F}}_{t}$. Since the modules in this phase are integral to both the encoding and decoding stages, we do not update them during the encoding process.
@@
Step 3. Contexture Feature Compression. Adhering to the methodologies established by current deep contextual coding frameworks (Li et al., 2022; Sheng et al., 2022; Li et al., 2023; Chen et al., 2023), we employ a multi-scale spatial-temporal Contexture Feature Encoder. This encoder generates the latent feature $\mathbf{Y}_{t}$ using the current frame $X_{t}$ and previously generated warped features $\bar{\mathbf{F}}_{t}$. Similar to Step 1, we again quantize $\mathbf{Y}_{t}$ into $\hat{\mathbf{Y}}_{t}$ and subsequently transmit it losslessly using arithmetic coding alongside a hyper-prior-based entropy model. Consequently, a Contexture Hyper-prior Feature Encoder is deployed to produce a hyper-prior feature. In this step, we enable the online updating of Contexture Feature Encoder and Contexture Hyper-prior Feature Encoder. For the complex module, Contexture Feature Encoder, parallel adaptors are integrated, whereas for the compact module, Contexture Hyper-prior Feature Encoder, serial adaptors are utilized.
@@
Step 4. Frame Reconstruction. Finally, we revert the quantized latent feature $\hat{\mathbf{Y}}_{t}$ back into the reconstructed frame $\hat{X}_{t}$ by utilizing a multi-scale spatial-temporal Contexture Feature Decoder, with assistance from the warped features $\bar{\mathbf{F}}_{t}$. This step is also used during the decoding process, hence there is no updating of any modules.
@@ 4. Experiments
Figure 4. Rate-distortion (i.e., bit-rates vs PSNR) performance comparison between our and other state-of-the-art video compression methods on the standard video compression benchmarks, UVG, MCL-JCV, HEVC Class B, C, D, and E.
@@ 4.1.1. Datasets
To evaluate the effectiveness of our approach, we implemented it on video sequences from six of the most representative video compression datasets: UVG (UVG, [n. d.]) and MCL-JCV (Wang et al., 2016), HEVC (Sullivan et al., 2012) Class B, C, D, and E. UVG and MCL-JCV are public collections of high-quality video sequences used for evaluating the compression algorithm, where all videos are $1920\times 1080$ pixels. HEVC datasets are selected by the standardization committees ISO/IEC and ITU-T for the assessment of new coding technologies. They encompass videos with resolutions ranging from 416$\times$240 to 1920$\times$1080 featuring diverse scenes. Furthermore, we assessed the adaptability of our method using the most widely used cardiac medical MRI sequence dataset, ACDC (Bernard et al., 2018), where we utilized 10 scans from Patient 101, each with a resolution of $256\times 232$. We measured the bit-rates by bits-per-pixel (bpp) and assessed the quality of compression by using the peak signal-to-noise ratio (PSNR).
@@ 4.1.2. Baseline Codecs
Our evaluation aimed to assess the efficacy of our method relative to both traditional video codecs and a range of NVC methods, including standard and content-adaptive variants. For traditional codecs, we utilized the H.265/HEVC (Sullivan et al., 2012) and H.266/VVC (Bross et al., 2021) standards, employing their reference software versions HM-16.20 (HM, [n. d.]) and VTM-11.2 (vtm, [n. d.]), respectively. Our analysis also included two leading NVCs as baselines: DCVC (Li et al., 2021) and its enhanced version, DCVC_DC (Li et al., 2023), with the latter representing the latest advancement in NVC study to the best of our knowledge. It surpasses traditional video codecs in performance, sharing the high-level syntax with DCVC but featuring a more intricate architecture. Additionally, we examined two content-adaptive NVC methods: one by Lu et al. (Lu et al., 2020) and another by Tang et al. (Tang et al., 2023) with Tang’s method being the most recent development in the field of content-adaptive NVC to the best of our knowledge. To ensure a fair comparison of videos, we recalibrated the performance metrics for DCVC, DCVC_DC, H.266, and H.265 using publicly available codebases and pre-trained models. Our setup involved an I-Frame interval of 32 and the encoding of 96 frames from each video, consistent with (Li et al., 2023). For each MRI scan from the ACDC dataset, we adopted an I-Frame interval of 30 and encoded 30 slices. For the other codecs, due to the unavailability of public implementations, we relied on performance data as reported in their original studies.